Grade 12th passAlgebra

Simplify combining like terms:
21b – 32 + 7b – 20b.....

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

When term have the same algebraic factors, they are like terms.
Then,
= (21b + 7b–20b)–32
= b (21 + 7–20)–32
= b (28 - 20)–32
= b (8) - 32
= 8b –32

Thanks